What is the threshold frequency that is required to emit photoelectron from selenium?


1
Expert's answer
2022-05-18T15:27:51-0400
"hf_{\rm th}=\phi"

"f_{\rm th}=\phi/h\\
={\rm(5.11\: eV\times 1.6*10^{-19}\: J/eV)/(6.62*10^{-34}\: J\cdot s)}\\
=1.24*10^{15}\:\rm Hz"